# Document-Level In-Context Few-Shot Relation Extraction via Pre-Trained Language Models

### v2

@@ 8 Discussion
Benefits: Our REPLM framework offers many benefits in practice: (1) REPLM eliminates the need for named entity recognition pipelines in our task and thus the noise that comes with it; (2) REPLM does not require human annotations but leverages in-context few-shot learning; and (3) REPLM offers great flexibility as it allows to incorporate new relations without re-training.
Implications: Our findings have direct implications for constructing datasets and benchmarking methods aimed at relation extraction from the textual documents. Evidently, earlier datasets such as DocRED (Yao et al., 2019), the largest dataset for document-level relation extraction publicly available, lack comprehensiveness and thus miss important – but correct – annotations. This may penalize correct methods and instead favor overly simple output during benchmarking. Hence, it will be important for future research to develop new datasets and effective evaluation paradigms.
Broader Impact Statement
This paper presents work whose primary goal is to advance the field of machine learning. We would like to emphasize the positive and negative implications that underlie machine learning more broadly. Our REPLM may help to fill the gaps in existing knowledge bases. It is likely that such gaps are particularly dominant for marginalized groups, so that our work can have a positive impact for improving coverage in knowledge bases for diverse groups. Still, LMs are far from perfect as their performance may vary. Hence, careful and responsible use is indicated, especially when filling in knowledge bases with important societal or ethical implications or sensitive content more broadly.
@@ Appendix A Related Work
In-context few-shot learning of LMs: In-context few-shot learning has been widely adopted in text classification tasks (Holtzman et al., 2021; Liu et al., 2022a; Lu et al., 2022a; Min et al., 2022a, b; Zhao et al., 2021) and further extended to other tasks such as question answering (Holtzman et al., 2021; Liu et al., 2022a; Min et al., 2022b), fact retrieval (Zhao et al., 2021), table-to-text generation (Liu et al., 2022a), and mapping utterances to meaning representations (Rubin et al., 2022). However, we are not aware of any earlier work that leveraged the in-context few-shot learning paradigm for document-level relation extraction.
LMs as knowledge bases: Research has focused on probing the knowledge in LMs. For example, Petroni et al. (2019) introduced the LAMA dataset, a dataset with cloze-style templates for different relations, which allows to probe factual knowledge in LMs. Many works have been introduced to achieve state-of-the-art results via prompt-tuning (Hao et al., 2022; Lester et al., 2021; Li and Liang, 2021; Liu et al., 2021b, 2022b; Newman et al., 2022; Perez et al., 2021; Poerner et al., 2020; Shin et al., 2020; Zhong et al., 2021). Yet, some works further find that, when evaluated as knowledge bases, LMs suffer from inconsistency (AlKhamissi et al., 2022; Elazar et al., 2021), learn shallow heuristics rather than facts (Elazar et al., 2022), have inferior performance in the long tail (Kandpal et al., 2022), and exhibit prompt bias (Cao et al., 2021).
@@
Relation extraction via pattern-based and statistical methods: A detailed review of the different methods is provided in Pawar et al. (2017) and Weikum et al. (2021), while we only present a brief summary here. Early works extracted relations from text via pattern-based extraction methods. Specifically, these works introduced automated methods to extract textual patterns corresponding to each relation and each entity type (Carlson et al., 2010; Jiang et al., 2017; Nakashole et al., 2012). Their main limitation is that the automatically constructed patterns involve many mistakes, which, in turn, require human experts to examine and correct them (Han et al., 2020).
Another research stream focused on relation extraction via statistical methods. Examples are crafting custom features for relation classification (Jiang and Zhai, 2007; Nguyen et al., 2007), designing customized kernels for support vector machines(Nguyen et al., 2007; Wang, 2008; Zhang et al., 2006b, a), graphical modeling and inference of relations (Sarawagi and Cohen, 2004; Yu and Lam, 2010), and leveraging knowledge graph embeddings for relation prediction (Lin et al., 2015; Wang et al., 2014).
@@
Relation extraction via neural networks: Initial methods for relation extraction based on neural network approaches made use of convolutional neural network (CNN) (Zeng et al., 2014) and long short-term memory (LSTM) (Zhou et al., 2016) architectures. These works process the pre-computed word embeddings and then classify the relation for the given named entity pair. Follow-up works proposed joint learning of entity extraction and relation classification, again via CNN (Adel and Schütze, 2017; Zheng et al., 2017) and LSTM (Katiyar and Cardie, 2017; Miwa and Bansal, 2016) architectures. However, these models are not flexible enough to model the complex interactions between named entities to classify the relation, as compared to LMs.
@@
Relation extraction via LMs: State-of-the-art methods for relation extraction are based on fine-tuning pre-trained LMs. Specifically, these methods use pre-trained LMs such as BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), and SciBERT (Beltagy et al., 2019) and fine-tuned them for relation extraction. For instance, Wang et al. (2019) fine-tuned BERT to classify the relation between each named entity pair in a given sentence. There have been various follow-up works to improve performance by learning complex dependency between named entities. To achieve this, Wang and Lu (2020) jointly trained LSTM and BERT to get two distinct representations of the entities; Zhang et al. (2021) further incorporate semantic segmentation module into the fine-tuning of BERT; Zhou et al. (2021) propose adaptive thresholding and localized context pooling; Xu et al. (2021) explicitly model the dependencies between entity mentions; Paolini et al. (2021) augmented the original sentences with the entity and relation types; Tan et al. (2022) use axial attention module for learning the interdependency among named entity pairs; Wang Xu and Zhao (2022) propose sentence importance estimation; Xiao et al. (2022) include additional tasks such as coreference resolution, entity typing, and evidence retrieval; Xu et al. (2023) improves the model performance via synthetic data generation; Hu et al. (2023) incorporates rationale extraction from the sentence; and Zhang et al. (2023a) leverages self-distillation to facilitate relational reasoning. However, all of these works require the named entities to be annotated and provided as input at both training and test time.
Appendix B DocRED Dataset
@@
For our experiments, we use DocRED (Yao et al., 2019), the largest publicly available dataset for document-level relation extraction. We provide the detailed statistics of each relation type in Tables 5 and 6 (note: the different columns compare the different subsets for distant supervision, human-annotated training, and human-annotated dev).
@@
Pre-processing. The original documents in the DocRED dataset are provided only in a tokenized format, e. g., the document is represented as a list of token, where each punctuation mark and word is a different token. We follow the earlier works (Cabot and Navigli, 2021; Yao et al., 2019) and concatenate the tokens with a white space in between to construct the entire document. This approach may introduce typos in the documents; for instance, the original text “Tarzan’s Hidden Jungle is a 1955 black-and-white film ...” is reconstructed as “Tarzan ’s Hidden Jungle is a 1955 black - and - white film ...”. We initially tried to fix these typos via spelling correction libraries, such as FastPunct [^12] https://pypi.org/project/fastpunct/, but later found that the typos are propagated to the labels, which may impede performance and eventually comparability of our results. Therefore, we decided to follow the same pre-processing as earlier works, as it allows us to operate on the same labels as in earlier work and thus ensures comparability of our results. Sec. G shows some examples of documents after the pre-processing step.
@@ Appendix C Details on Sentence-Level Relation Extraction
CONLL04 (Roth and Yih, 2004) is consisting of sentences collected from the news articles. The authors manually annotated the entities and 5 relation types for each sentence. Following the earlier literature, we used the same splits as Eberts and Ulges (2020). The detailed statistics for each relation type are given in Table 7.
@@
ADE (Gurulingappa et al., 2012) contains the sentences from biomedical domain and it has only one relation type, which is adverse effect. The original dataset contains 10 folds of train and test splits. Following the earlier work (Cabot and Navigli, 2021), we use the test set of the first fold for the evaluation. The statistics are given in Table 9.
@@
We perform an ablation study to observe the effect of using distantly-supervised vs. human-annotated documents as in-context few-shot examples. We perform such comparison on our four framework variants, namely REPLM (random fixed), REPLM (random all), REPLM (best $\ominus$), and REPLM (best $\oplus$). For methods with random in-context examples, the performance may be subject to variability across which seed is picked (whereas the performance is deterministic for the other methods), and, hence, we report the standard deviation for this subset of the methods by averaging the performance across 10 runs. Of note, to directly compare the impact of in-context examples, we deliberately considered our REPLM variants without aggregation over the multiples sets of in-context examples.
Tables 10 to 21 show the comparison between distant supervision vs. human annotation. For random in-context examples (i. e., REPLM (random fixed) and REPLM (random all)), distant supervision and human annotation performs at the same level. For retrieving the semantically most similar context examples (i. e., REPLM (best $\ominus$) and REPLM (best $\oplus$)), we observe cases where distant supervision actually improves the result (e. g., P6, P155, P179). However, the overall performance is largely similar. This confirms our choice of using distantly-supervised documents as in-context examples and eliminates the need for human annotation.
@@ Appendix E Implementation Details
In our REPLM framework, we retrieve the semantically most-relevant in-context examples for each dev document. For this, we encode the documents via SBERT (Reimers and Gurevych, 2019) to calculate the embeddings and retrieve the most-relevant documents based on the cosine-similarity of the embeddings. We use the following fixed parameters in our framework (if not specified otherwise): $N=20$, $K=5$ $L=5$, $\tau=0.1$, and $\theta=0.2$. As the sentences in sentence-level relation extraction datasets (CONLL04, NYT, and ADE) are shorter than the documents in DocRED, we used more in-context examples for these datasets, which is $K=11$. The other parameters are the same as before.
We run all our experiments on NVIDIA Tesla V100-SXM2 32GB with a batch size 4. For DocRED, on average, each batch is processed in $\sim$17.80 seconds. As a result, the dev set (998 documents) is processed in $\sim$74.17 minutes for each relation type.
@@ Appendix H Full Results on Overall Performance
Tables 22 to 33 provide the detailed results of relation extraction of all methods on all relation types. For methods with random in-context examples, the performance may be subject to variability across which seed is picked (whereas the performance is deterministic for the other methods), and, hence, we report the standard deviation for this subset of the methods by averaging the performance across 10 runs. The evaluation is done based on the human annotations of the dev documents.
@@ Appendix I Full Results on Comparison Against External Knowledge
Tables 34 to 41 provide the detailed results of relation extraction of all methods on all relation types. The evaluation is done based on checking the correctness of extracted relations against both the human annotations of the dev documents and an external knowledge base.
Overall, we see that our our REPLM and REPLM (params adj) outperforms REBEL in most relations, when further including the relations from an external knowledge base. The performance improvement becomes more striking for the relation types having a large number of knowledge triplets. For instance, for the relation P17 (country), our REPLM achieves an F1 score of 56.14, whereas REBEL-sent can achieve less than the half of the performance, F1 score of 21.17. We want to highlight that, these methods performed at similar level when compared against only human-annotations as ground-truth (REPLM with 27.74 F1 score vs. REBEL-sent with 25.73 F1 score, see Sec. H). Therefore, it shows the importance of evaluating against an external knowledge base. Even a larger performance gap can be found in relation P27 (country of citizenship), our REPLM and REPLM (params adj) achieve the F1 scores 39.52 and 46.15, respectively, whereas REBEL and REBEL-sent can achieve only F1 scores 3.47 and 8.24. These results further confirm the effectiveness of our REPLM framework, agreeing with the results from the earlier sections.
@@
In our original REPLM framework, we used a fixed temperature $\tau$ and a fixed probability threshold $\theta$ as our REPLM does not require a human-annotated training documents to tune the hyperparameters. Tables 42 to 47 show the sensitivity of REPLM on various hyperparameter selections for all relation types individually. Overall, we observe that our REPLM framework is robust to different hyperparameter: the performance remains at the same level for most variations. (Only exception is $\theta=0.5$, where such high probability threshold results in discarding the correctly extracted relations. This leads to much lower recall, and therefore, F1 score.) This also explains why the performance gap between REPLM and REPLM (params adj) is small.
@@
We performed an ablation study to show how the performance of REPLM framework changes with different number of in-context examples (i. e., varying $K$). For this, we run the same experiment on CONLL04 dataset with $K=3,5,7,9,11$ while keeping all the other parameters fixed. Table 49 shows the performance on each relation type and the overall performance. We observe that more in-context examples yield better F1 score in general. Informed by this observation, we used the highest number of in-context examples ($K$) in our experiments. For the document-level relation extraction dataset DocRED, we use $K=5$ as more than 5 documents do not fit into the context window of our REPLM framework.
Table 49: Comparing the performance across varying number of in-context examples $K$ for the dataset CONLL04. Shown are F1 scores on each relation and overall (Micro) F1 score.
@@
We test whether our REPLM framework is actually learning to extract the relation from the document or simply retrieving the facts from its own memory in the specified format. For this, we re-construct the sentences from CONLL04 (Roth and Yih, 2004) with fake entity names that are not mentioned anywhere in the web. As a result, we push the limits of our REPLM and test whether the model correctly extracts the relations about the entities that appear only in the input sentence. To the best of our knowledge, we are the first to design such an experiment to shed light on the learning abilities of a LM in the context of relation extraction.
Table 50 compares the performance of our REPLM based on the original dataset vs. the same dataset with fake entities. There is only a slight decrease in the overall performance with the fake entities (F1 score of 70.47 vs. 72.94), confirming that our REPLM is actually learning to extract the relations from the context. After this confirmation, we further investigate “how” our REPLM learns to extract relations, by contrasting its behavior against the presence of adversarial in-context examples (see Appendix N).
We conjecture that the slight decrease in the performance can be attributed to mainly two factors: (1) In some cases, the relation between the entities becomes unclear without knowing what these entities actually are. As can be seen from the in-context examples at Sec. M.1, it is hard to infer “work for” relation between “Entity62” and “Entity22” from “ normal-$\ldots$ , said Lt. Entity62 of the Entity22”, which mitigates the model performance. (2) The random entity names (e. g., “Entity10”, “Entity55” etc.) are not seen during the pre-training of the model. Therefore, the likelihood of generating the fake entities from the context is not the same as generating the real names that appeared during the pre-training, which can impact the performance of a LM.
Table 50: Comparing the performance on the original dataset vs. the dataset with random entity names. Shown are F1 scores on each relation and overall (Micro) F1 score.
@@
Table 51 shows the overall results. Furthermore, Tables 52 to 57 reports the performance across each relation type. The performance clearly drops as a result of including the adversarial in-context examples. This has two crucial implications: (1) The quality of the labels is important. Overall, we perform our experiments with distantly-supervised documents that are automatically annotated with certain heuristics to ensure quality (as explained in the main paper). If these are ignored, the resulting in-context examples can mislead the LM. (2) Our framework actually learns from the in-context examples. When it learns the relations with noisy labels, this also reflects into its performance. Further, comparing the performance between “100 % clean” and “50 % clean”, we see that the relation extractions stays at the same level. This informs us that, if some adversarial documents exist in the dataset, filtering them out is a viable option and there is no need to find another method to fix their labels.

### v4

@@ 9 Discussion
Benefits: Our REPLM framework offers many benefits in practice: (1) REPLM eliminates the need for named entity recognition pipelines in our task and thus the error propagated with it; (2) REPLM does not require human annotations but leverages in-context few-shot learning; and (3) REPLM offers great flexibility as it allows to incorporate new relations and new backbone LMs without re-training. Our study further identifies earlier datasets, such as DocRED (Yao et al., 2019), lack comprehensiveness, and thus miss important – but correct – annotations. This may penalize correct methods during benchmarking, suggesting the need of more effective evaluation paradigms.
Broader Impact: Our REPLM can help bridge gaps in knowledge bases, particularly for marginalized groups, improving coverage for diverse populations. However, as LM performance can vary, careful and responsible use is necessary when addressing societal, ethical, or sensitive content.
@@ Appendix A Related Work
In-context few-shot learning of LMs: In-context few-shot learning has been widely adopted in text classification tasks (Holtzman et al., 2021, Liu et al., 2022a, Lu et al., 2022a, Min et al., 2022a; b, Zhao et al., 2021b) and further extended to other tasks such as question answering (Holtzman et al., 2021, Liu et al., 2022a, Min et al., 2022b), fact retrieval (Zhao et al., 2021b), table-to-text generation (Liu et al., 2022a), and mapping utterances to meaning representations (Rubin et al., 2022). However, we are not aware of any earlier work that leveraged the in-context few-shot learning paradigm for document-level relation extraction.
LMs as knowledge bases: Research has focused on probing the knowledge in LMs. For example, Petroni et al. (2019) introduced the LAMA dataset, a dataset with cloze-style templates for different relations, which allows to probe factual knowledge in LMs. Many works have been introduced to achieve state-of-the-art results via prompt-tuning (Hao et al., 2022, Lester et al., 2021, Li & Liang, 2021, Liu et al., 2021b; 2022b, Newman et al., 2022, Perez et al., 2021, Poerner et al., 2020, Shin et al., 2020, Zhong et al., 2021). Yet, some works further find that, when evaluated as knowledge bases, LMs suffer from inconsistency (AlKhamissi et al., 2022, Elazar et al., 2021), learn shallow heuristics rather than facts (Elazar et al., 2022), have inferior performance in the long tail (Kandpal et al., 2022), and exhibit prompt bias (Cao et al., 2021).
@@
Relation extraction via pattern-based and statistical methods: A detailed review of the different methods is provided in Pawar et al. (2017) and Weikum et al. (2021), while we only present a brief summary here. Early works extracted relations from text via pattern-based extraction methods. Specifically, these works introduced automated methods to extract textual patterns corresponding to each relation and each entity type (Carlson et al., 2010, Jiang et al., 2017, Nakashole et al., 2012). Their main limitation is that the automatically constructed patterns involve many mistakes, which, in turn, require human experts to examine and correct them (Han et al., 2020).
Another research stream focused on relation extraction via statistical methods. Examples are crafting custom features for relation classification (Jiang & Zhai, 2007, Nguyen et al., 2007), designing customized kernels for support vector machines(Nguyen et al., 2007, Wang, 2008, Zhang et al., 2006a; b), graphical modeling and inference of relations (Sarawagi & Cohen, 2004, Yu & Lam, 2010), and leveraging knowledge graph embeddings for relation prediction (Lin et al., 2015, Wang et al., 2014).
@@
Relation extraction via neural networks: Initial methods for relation extraction based on neural network approaches made use of convolutional neural network (CNN) (Zeng et al., 2014) and long short-term memory (LSTM) (Zhou et al., 2016) architectures. These works process the pre-computed word embeddings and then classify the relation for the given named entity pair. Follow-up works proposed joint learning of entity extraction and relation classification, again via CNN (Adel & Schütze, 2017, Zheng et al., 2017) and LSTM (Katiyar & Cardie, 2017, Miwa & Bansal, 2016) architectures. However, these models are not flexible enough to model the complex interactions between named entities to classify the relation, as compared to LMs.
@@
Relation extraction via LMs: State-of-the-art methods for relation extraction are based on fine-tuning pre-trained LMs. Specifically, these methods use pre-trained LMs such as BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), and SciBERT (Beltagy et al., 2019) and fine-tuned them for relation extraction. For instance, Wang et al. (2019) fine-tuned BERT to classify the relation between each named entity pair in a given sentence. There have been various follow-up works to improve performance by learning complex dependency between named entities. To achieve this, Wang & Lu (2020) jointly trained LSTM and BERT to get two distinct representations of the entities; Zhang et al. (2021) further incorporate semantic segmentation module into the fine-tuning of BERT; Zhou et al. (2021) propose adaptive thresholding and localized context pooling; Xu et al. (2021a) explicitly model the dependencies between entity mentions; Paolini et al. (2021) augmented the original sentences with the entity and relation types; Tan et al. (2022) use axial attention module for learning the interdependency among named entity pairs; Wang Xu & Zhao (2022) propose sentence importance estimation; Xiao et al. (2022) include additional tasks such as coreference resolution, entity typing, and evidence retrieval; Xu et al. (2023) improves the model performance via synthetic data generation; Hu et al. (2023) incorporates rationale extraction from the sentence; and Zhang et al. (2023a) leverages self-distillation to facilitate relational reasoning. However, all of these works require the named entities to be annotated and provided as input at both training and test time.
Appendix B Details on Document-level Relation Extraction Datasets
B.1 DocRED
@@
For our experiments, we mainly use DocRED (Yao et al., 2019), the largest publicly available dataset for document-level relation extraction. We provide the detailed statistics of each relation type in Tables 6 and 7 (note: the different columns compare the different subsets for distant supervision, human-annotated training, and human-annotated dev).
@@
Pre-processing. The original documents in the DocRED dataset are provided only in a tokenized format, e. g., the document is represented as a list of token, where each punctuation mark and word is a different token. We follow the earlier works (Cabot & Navigli, 2021, Yao et al., 2019) and concatenate the tokens with a white space in between to construct the entire document. This approach may introduce typos in the documents; for instance, the original text “Tarzan’s Hidden Jungle is a 1955 black-and-white film ...” is reconstructed as “Tarzan ’s Hidden Jungle is a 1955 black - and - white film ...”. We initially tried to fix these typos via spelling correction libraries, such as FastPunct [^12] https://pypi.org/project/fastpunct/, but later found that the typos are propagated to the labels, which may impede performance and eventually comparability of our results. Therefore, we decided to follow the same pre-processing as earlier works, as it allows us to operate on the same labels as in earlier work and thus ensures comparability of our results. Sec. G shows some examples of documents after the pre-processing step.
@@
B.2 CDR and GDA
CDR (Li et al., 2016b) contains the abstracts from PubMed (https://pubmed.ncbi.nlm.nih.gov/) but it contains only one relation type, which is the chemical-induced disease. We use the original splits in our work. The dataset statistics can be found in Table 8.
Table 8: CDR statistics
Relation Name # Docs in Train # Relations in Train # Docs in Validation # Relations in Validation # Docs in Test # Relations in Test
Chemical-Induced Disease 500 1038 500 1012 500 1066
GDA (Wu et al., 2019) is another dataset that offers a collection of documents from the medical domain. The documents are again the abstracts of PubMed. The dataset contains only one relation type, which is gene-disease association. We use the original test split for the evaluation of our work. The statistics of GDA can be found in Table 9.
Table 9: GDA statistics
Relation Name # Docs in Train # Relations in Train # Docs in Validation # Relations in Validation # Docs in Test # Relations in Test
Gene-Disease Association 29192 44841 _ _ 1000 1502
@@ Appendix C Details on Sentence-Level Relation Extraction
CONLL04 (Roth & Yih, 2004) is consisting of sentences collected from the news articles. The authors manually annotated the entities and 5 relation types for each sentence. Following the earlier literature, we used the same splits as Eberts & Ulges (2020). The detailed statistics for each relation type are given in Table 10.
@@
ADE (Gurulingappa et al., 2012) contains the sentences from biomedical domain and it has only one relation type, which is adverse effect. The original dataset contains 10 folds of train and test splits. Following the earlier work (Cabot & Navigli, 2021), we use the test set of the first fold for the evaluation. The statistics are given in Table 12.
@@
We perform an ablation study to compare the effect of using distantly-supervised vs. human-annotated documents as in-context few-shot examples. We perform such comparison using the four variants of our framework variants, i.e., REPLM (random fixed), REPLM (random all), REPLM (best $\ominus$), and REPLM (best $\oplus$). For methods with random in-context examples, the performance may be subject to variability across which seed is picked (whereas the performance is deterministic for the other methods), and, hence, we report the standard deviation for this subset of the methods by averaging the performance across 10 runs. Of note, to directly compare the impact of in-context examples, we deliberately considered our REPLM variants without aggregation over the multiples sets of in-context examples.
Tables 13 to 24 show the comparison between distant supervision vs. human annotation. For random in-context examples (i. e., REPLM (random fixed) and REPLM (random all)), distant supervision and human annotation performs at the same level. For retrieving the semantically most similar context examples (i. e., REPLM (best $\ominus$) and REPLM (best $\oplus$)), we observe cases where distant supervision actually improves the result (e. g., P6, P155, P179). However, the overall performance is largely similar. This confirms our choice of using distantly-supervised documents as in-context examples and eliminates the need for human annotation.
@@ Appendix E Implementation Details
In our REPLM framework, we retrieve the semantically most-relevant in-context examples for each dev document. For this, we encode the documents via SBERT (Reimers & Gurevych, 2019) to calculate the embeddings and retrieve the most-relevant documents based on the cosine-similarity of the embeddings. We use the following fixed parameters in our framework (if not specified otherwise): $N=20$, $K=5$ $L=5$, $\tau=0.1$, and $\theta=0.2$. As the sentences in sentence-level relation extraction datasets (CONLL04, NYT, and ADE) are shorter than the documents in DocRED, we used more in-context examples for these datasets, which is $K=11$. The other parameters are the same as before.
We run all of the experiments of REPLM on NVIDIA Tesla V100-SXM2 32GB with a batch size of 4. For DocRED, on average, each batch is processed in $\sim$17.80 seconds. As a result, the dev set (998 documents) is processed in $\sim$74.17 minutes for each relation type.
For REPLM variants, we use gpt-3.5-turbo and gpt-4o of OpenAI and Llama-3.1-8B and Llama-3.1-70B from Meta as our backbone LMs. Similar to our design choice with GPT-JT, we opt for deterministic outputs from these LM backbones, which is done by choosing a low temperature such as 0.001. If not specified otherwise, we use exactly the same parameter configuration with our REPLM framework, specifically $N=20$, $K=5$ $L=5$, $\tau=0.1$, and $\theta=0.2$.
@@ Appendix H Full Results on Overall Performance
Tables 25 to 36 provide the detailed results of relation extraction of all methods on all relation types. For methods with random in-context examples, the performance may be subject to variability across which seed is picked (whereas the performance is deterministic for the other methods), and, hence, we report the standard deviation for this subset of the methods by averaging the performance across 10 runs. The evaluation is done based on the human annotations of the dev documents.
@@ Appendix I Full Results on Comparison Against External Knowledge
Tables 37 to 44 provide the detailed results of relation extraction of all methods on all relation types. The evaluation is done based on checking the correctness of extracted relations against both the human annotations of the dev documents and an external knowledge base.
Overall, we see that our REPLM and REPLM (params adj) outperform REBEL in most relations, when further including the relations from an external knowledge base. The performance improvement becomes more striking for the relation types having a large number of knowledge triplets. For instance, for the relation P17 (country), our REPLM achieves an F1 score of 56.14, whereas REBEL-sent can achieve less than half of the performance, that is, an F1 score of 21.17. We want to highlight that, these methods performed at a similar level when compared against only human-annotations as ground-truth (REPLM with 27.74 F1 score vs. REBEL-sent with 25.73 F1 score, see Sec. H). Therefore, it shows the importance of evaluating against an external knowledge base. Even a larger performance gap can be found in relation P27 (country of citizenship), our REPLM and REPLM (params adj) achieve the F1 scores 39.52 and 46.15, respectively, whereas REBEL and REBEL-sent can achieve only F1 scores 3.47 and 8.24. These results further confirm the effectiveness of our REPLM framework, agreeing with the results from the earlier sections.
@@
In our original REPLM framework, we used a fixed temperature $\tau$ and a fixed probability threshold $\theta$ as our REPLM does not require a human-annotated training documents to tune the hyperparameters. Tables 45 to 50 show the sensitivity of REPLM on various hyperparameter selections for all relation types individually. Overall, we observe that our REPLM framework is robust to different hyperparameter: the performance remains at the same level for most variations. (Only exception is $\theta=0.5$, where such high probability threshold results in discarding the correctly extracted relations. This leads to much lower recall, and therefore, F1 score.) This also explains why the performance gap between REPLM and REPLM (params adj) is small.
@@
We performed an ablation study to show how the performance of REPLM framework changes with different number of in-context examples (i. e., varying $K$). For this, we run the same experiment on CONLL04 dataset with $K=3,5,7,9,11$ while keeping all the other parameters fixed. Table 52 and Table 53 show the performance of REPLM and REPLM+GPT3.5, respectively, for each relation type and the overall performance. We observe that more in-context examples yield better F1 scores in general. Informed by this observation, we used the highest number of in-context examples ($K$) in our experiments. For the document-level relation extraction dataset DocRED, we use $K=5$ as more than 5 documents do not fit into the context window of our REPLM framework.
Table 52: Comparing the performance of REPLM across varying number of in-context examples $K$ for the dataset CONLL04. Shown are F1 scores on each relation and overall (Micro) F1 score.
@@
Table 53: Comparing the performance of REPLM+GPT3.5 across varying number of in-context examples $K$ for the dataset CONLL04. Shown are F1 scores on each relation and overall (Micro) F1 score.
3 91.63 74.45 71.08 82.41 77.01 77.94
5 93.85 75.58 69.81 81.29 79.56 78.60
7 91.93 75.78 73.00 80.52 79.85 78.81
9 94.92 76.36 72.45 81.11 80.12 79.49
11 93.75 76.06 73.32 84.31 80.98 80.19
@@
We test whether our REPLM framework is actually learning to extract the relation from the document or simply retrieving the facts from its own memory in the specified format. For this, we re-construct the sentences from CONLL04 (Roth & Yih, 2004) with fake entity names that are not mentioned anywhere in the web. As a result, we push the limits of our REPLM and test whether the model correctly extracts the relations about the entities that appear only in the input sentence. To the best of our knowledge, we are the first to design such an experiment to shed light on the learning abilities of a LM in the context of relation extraction.
Table 54 compares the performance of our REPLM based on the original dataset vs. the same dataset with fake entities. There is only a slight decrease in the overall performance with the fake entities (F1 score of 70.47 vs. 72.94), confirming that our REPLM is actually learning to extract the relations from the context. Further, Table 55 shows that our findings from REPLM are transferable to REPLM+GPT3.5, i.e., to more advanced LMs. After this confirmation, we further investigate “how” our REPLM learns to extract relations, by contrasting its behavior against the presence of adversarial in-context examples (see Appendix N).
We conjecture that the slight decrease in the performance can be attributed to mainly two factors: (1) In some cases, the relation between the entities becomes unclear without knowing what these entities actually are. As can be seen from the in-context examples at Sec. M.1, it is hard to infer “work for” relation between “Entity62” and “Entity22” from “ $\ldots$ , said Lt. Entity62 of the Entity22”, which mitigates the model performance. (2) The random entity names (e. g., “Entity10”, “Entity55” etc.) are not seen during the pre-training of the model. Therefore, the likelihood of generating the fake entities from the context is not the same as generating the real names that appeared during the pre-training, which can impact the performance of a LM.
Table 54: Comparing the performance of REPLM on the original dataset vs. the dataset with random entity names. Shown are F1 scores on each relation and overall (Micro) F1 score.
@@
Table 55: Comparing the performance of REPLM+GPT3.5 on the original dataset vs. the dataset with random entity names. Shown are F1 scores on each relation and overall (Micro) F1 score.
Dataset Kill Live_In Located_In OrgBased_In Work_For Overall
Original 93.75 76.06 72.32 84.31 80.98 80.19
Random Entities 88.89 74.65 68.93 85.59 85.71 79.65
@@
Table 56 shows the overall results. Furthermore, Tables 57 to 62 reports the performance across each relation type. The performance clearly drops as a result of including the adversarial in-context examples. This has two crucial implications: (1) The quality of the labels is important. Overall, we perform our experiments with distantly-supervised documents that are automatically annotated with certain heuristics to ensure quality (as explained in the main paper). If these are ignored, the resulting in-context examples can mislead the LM. (2) Our framework actually learns from the in-context examples. When it learns the relations with noisy labels, this also reflects into its performance. Further, comparing the performance between “100 % clean” and “50 % clean”, we see that the relation extractions stays at the same level. This informs us that, if some adversarial documents exist in the dataset, filtering them out is a viable option and there is no need to find another method to fix their labels.
@@
Appendix O Detailed Performance Comparison against Other In-Context Learning Methods
In this section, we compare the performance of our REPLM against the in-context learning methods developed for sentence-level relation extraction task (Wadhwa et al., 2023, Wan et al., 2023). As explained in Sec. 2, these models are not scalable to document-level. Further, at the time of writing, the implementation of (Wadhwa et al., 2023) is not publicly available, therefore, we compare our REPLM against GPT-RE (Wan et al., 2023).
Figure 5: Comparing the performance of our REPLM against GPT-RE (Wan et al., 2023) on CONLL04 dataset.
Figure 5 shows the performance of our REPLM and GPT-RE (Wan et al., 2023) for each relation, along with the overall performance (Micro F1). Our REPLM consistently outperforms GPT-RE on each relation and it achieves roughly 60% F1 score improvement over GPT-RE (72.94 vs. 45.78). We attribute the inferior performance of GPT-RE to mainly two reasons: (1) GPT-RE introduces the in-context examples of all relations into the same context, therefore, the model is forced to learn the classification of all relation types at one inference. With more relation types, the task becomes more difficult. (2) In most cases, GPT-RE outputs one of the relation types, even when there is no relation between the given entity pairs. Although the in-context examples include “no relation” instances, it is not possible to cover all variations of “no relation” cases.
[…]
Figure 6: Comparing the performance of our REPLM+GPT3.5 against GPT-RE (Wan et al., 2023) on CONLL04 dataset.
To better demonstrate the strength of our framework, we further compared GPT-RE against our REPLM+GPT3.5, in which the backbone LM is replaced by GPT-3.5, so that we use the same GPT model that GPT-RE leverages. Figure 6 shows that our REPLM+GPT3.5 outperforms GPT-RE by even more larger margins. Specifically, it achieves more than 75% F1 score improvement over GPT-RE (80.19 vs. 45.78). Further, our REPLM+GPT3.5 requires much less API calls than GPT-RE, therefore, it costs around 15 USD in comparison to 100 USD spending required from GPT-RE.